### v1

# Explicit Runge-Kutta Methods with MQ and IMQ-Radial Basis Functions

Abstract
This article presents a class of explicit Runge-Kutta methods with multiquadric (MQ) and inverse multiquadric (IMQ) radial basis functions (RBFs) to improve the accuracy of time integration for ordinary differential equations. By introducing RBF-based corrections derived from Taylor series expansions and optimally selecting the shape parameter, the method achieves a one-order increase in accuracy without additional stages. Convergence and stability analyses support the theoretical claims, and numerical experiments in MATLAB confirm the predicted performance.
keywords:
Initial value problem , Runge-Kutta methods , Shape parameter , Order of accuracy , Radial basis functions , Stability.
AMS Classification : 65L06
1 Introduction
Numerical methods for solving initial value problems (IVPs) of the form
$u^{\prime}=f(t,u),\quad u(t_{0})=u_{0},\quad a\leq t\leq b,$ (1.1)
where $u(t)$to $C^{\infty}[a,b]$ and $f(t,u)$ is class of $C^{\infty}$ has enormous use in physics, biology, engineering, and control theory. Among the many available techniques, Runge–Kutta (RK) methods are widely used due to their explicit structure, ease of implementation, and high-order accuracy. These methods compute a weighted average of intermediate slope evaluations between $(t_{n},u_{n})$ and $(t_{n+1},u_{n+1})$, and a classical explicit $r$-stage RK method attains order $r$ for $r\leq 4$ [1]. However, for $r\geq 5$, there exists an order barrier, and increasing the number of stages does not necessarily yield higher accuracy if polynomial basis is being used.
To address this limitation, Gu et al. [2] proposed Runge–Kutta methods enhanced with Gaussian radial basis functions (RBFs). Radial basis functions (RBFs) have been integrated into RK methods to enhance their accuracy without increasing the number of stages. This RBF-based method extends the classical approach by embedding a free parameter in the structure of each stage, inspired by the Gaussian RBF Euler method [3]. The idea is to reduce the local truncation error by choosing a free parameter such that the leading error term vanishes, thus improving the order of accuracy by one. The value of the free parameter or the so-called shape parameter demands the local smoothness of the solution.
The shape parameter $\epsilon\in\mathbb{R}\cup i\mathbb{R}$ plays a crucial role in tuning the method both the accuracy and stability of the methods. While there is no universal rule for determining its optimal value, various approaches including stability and accuracy have been used in the literature [2, 3, 4, 5, 6, 7] have even proposed using multiple shape parameters to further enhance accuracy. The use of shape parameters to suppress numerical oscillations or errors is not limited to RBF-based ODE solvers. Similar ideas appear in error inhibiting schemes [8] , Essentially Non-Oscillatory (ENO) and Weighted Essentially Non-Oscillatory (WENO) [9, 10] methods for solving hyperbolic PDEs, where the reconstruction uses adaptive criteria to control local behavior of the solution.
In this work, we adopt the idea of directly eliminating leading truncation error terms through the clever selection of shape parameters. In the literature [4, 5, 6], there are methods in which the optimal value of the shape parameter is used. Here, we analytically determine the shape parameters to cancel terms in the local truncation error expansion up to the desired order. This provides a straightforward and systematic mechanism for constructing high-order RBF-based methods. We extend this framework beyond Gaussian RBFs  [2] and analyze Runge–Kutta methods based on multiquadric (MQ-RBF) [11] and inverse multiquadric (IMQ-RBF) Euler methods [7]. Using the functional form of the RBF-modified Euler methods, we derive the corresponding multistage RK methods and provide local truncation error analysis to establish the convergence of the method. Note that the polynomial-based Runge-Kutta method is a special case of the RBF-based Runge-Kutta method. So, the RBF based Runge-Kutta method shows at least the same order of convergence as that of the classical Runge-Kutta method.
Throughout the paper, $u_{n}$ and $v_{n}$ denote the exact and numerical solutions of the IVP (1.1), respectively. Our analysis focuses on improving the convergence behavior and efficiency of these RBF-RK methods while maintaining explicitness and ease of implementation.
The manuscript is divided into the following sections. Section 2 deals with r-stage classical Runge Kutta method with order r, $r\leq 3$. Section 3 defines the MQ-RBF Runge-Kutta method and the IMQ-RBF Runge-Kutta method and derives the respective expression for the truncation error. In Section 4, we examine the convergence analysis of the MQ and IMQ Runge-Kutta methods. Numerical examples are performed in Section 5, and we demonstrate conclusion and insights in section 6.
2 Explicit Runge-Kutta method
The $r$-stage Runge-Kutta methods with step length $h$ can be written as,
$v_{n+1}=v_{n}+h\sum_{i=1}^{r}b_{i}k_{i},$ (2.1)
where,
$k_{1}$ $=f(t_{n},v_{n}),$
$k_{2}$ $=f(t_{n}+c_{2}h,v_{n}+ha_{21}k_{1}),$
$k_{3}$ $=f(t_{n}+c_{3}h,v_{n}+h(a_{31}k_{1}+a_{32}k_{2})),$
[…]
$k_{r}$ $=f\left(t_{n}+c_{r}h,v_{n}+h\sum_{j=1}^{r-1}a_{rj}k_{j}\right).$
Explicit r-stage Runge-Kutta methods in Butcher tableau form can be written with Butcher tableau representation as [1]
$c_{1}=0$
$c_{2}$ $a_{21}$
$c_{3}$ $a_{31}$ $a_{32}$
$\vdots$ $\vdots$ $\vdots$ $\ddots$
$c_{r}$ $a_{r1}$ $a_{r2}$dots $a_{r(r-1)}$
$b_{1}$ $b_{2}$dots $b_{r}$
Expanding (2.1) with Taylor’s series and equating the coefficients of h from both sides, we get,
$\sum_{j=1}^{r}{b_{j}}=1,$ (2.2)
$\sum_{j=1}^{i-1}{a_{ij}}=c_{i},i=1,...,r.$ (2.3)
2.1 Two-stage second-order methods
The two-stage Runge-Kutta method can be written as
$v_{n+1}$ $=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}),$ (2.4)
$k_{1}$ $=f(t_{n},v_{n}),$
$k_{2}$ $=f(t_{n}+c_{2}h,v_{n}+ha_{21}k_{1}),$
and the corresponding Butcher tableau is
$\begin{array}[]{c|cc}0&\\ c_{2}&a_{21}\\ \hline\cr&b_{1}&b_{2}\\ \end{array}$
Using the Taylor series expansion, the Runge-Kutta method (2.4) become second-order accurate if
$b_{2}c_{2}=\frac{1}{2},$ (2.5)
along with the conditions (2.2) and (2.3) converts two-stage Runge-Kutta methods to a one-parameter family.
$v^{(1)}$ $=v_{n}+c_{2}hf(t_{n},v_{n}),$
$v_{n+1}$ $=\left(1-\frac{1}{c_{2}}\right)\left(1-\frac{1}{2c_{2}}\right)v_{n}+\frac{1}{c_{2}}\left(1-\frac{1}{2c_{2}}\right)v^{(1)}+\frac{1}{2c_{2}}hf\left(t_{n}+c_{2}h,v^{(1)}\right)$
Butcher tableau representation of the same is
$\begin{array}[]{c|cc}0&\\ c_{2}&c_{2}\\ \hline\cr&1-\frac{1}{2c_{2}}&\frac{1}{2c_{2}}\\ \end{array}$
when $c_{2}\neq 0.$
2.2 Three-stage third-order methods
Three-stage Runge-Kutta method is of the form
$v_{n+1}$ $=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}+b_{3}k_{3}),$ (2.6)
$k_{1}$ $=f(t_{n},v_{n}),$
$k_{2}$ $=f(t_{n}+c_{2}h,v_{n}+ha_{21}k_{1}),$
$k_{3}$ $=f(t_{n}+c_{3}h,v_{n}+h(a_{31}k_{1}+a_{32}k_{2})).$
Butcher tableau representation of the method is given by
$\begin{array}[]{c|ccc}0&\\ c_{2}&a_{21}\\ c_{3}&a_{31}&a_{32}\\ \hline\cr&b_{1}&b_{2}&b_{3}\\ \end{array}$
Another form of three-stage Runge-Kutta method is
$v^{(1)}$ $=v_{n}+c_{2}hf(t_{n},v_{n}),$
$v^{(2)}$ $=\left(1-\frac{a_{31}}{c_{2}}\right)v_{n}+\frac{a_{31}}{c_{2}}v^{(1)}+a_{32}hf\left(t_{n}+c_{2}h,v^{(1)}\right),$
$v_{n+1}$ $=\left(1-\frac{1}{c_{2}}\left(b_{1}-\frac{a_{31}}{a_{32}}b_{2}\right)-\frac{b_{2}}{a_{32}}\right)v_{n}+\frac{1}{c_{2}}\left(b_{1}-\frac{a_{31}}{a_{32}}b_{2}\right)v^{(1)}+\frac{b_{2}}{a_{32}}v^{(2)}+b_{3}hf\left(t_{n}+c_{3}h,v^{(2)}\right).$
The method will be third-order accurate if,
$b_{2}c_{2}+b_{3}c_{3}$ $=\frac{1}{2},$ (2.7)
$b_{2}c_{2}^{2}+b_{3}c_{3}^{2}$ $=\frac{1}{3},$
$a_{32}b_{3}c_{2}$ $=\frac{1}{6},$
along with the conditions (2.2) and (2.3). Based on parameter value the families of three-stage third-order Runge-Kutta methods are given below.
I. $c_{2}\neq 0$, $c_{2}\neq\frac{2}{3}$, $c_{3}\neq 0$, $c_{2}\neq c_{3}$
$\begin{array}[]{c|ccc}0&\\ c_{2}&c_{2}\\ c_{3}&\frac{c_{3}(3c_{3}-3c_{2}^{2}-c_{3})}{c_{2}(2-3c_{2})}&\frac{c_{3}(c_{3}-c_{2})}{c_{2}(2-3c_{2})}\\ \hline\cr&\frac{-3c_{3}+6c_{2}c_{3}+2-3c_{3}}{6c_{2}c_{3}}&\frac{3c_{3}-2}{6c_{2}(c_{3}-c_{2})}&\frac{2-3c_{2}}{6c_{3}(c_{3}-c_{2})}\\ \end{array}$
II. $b_{3}\neq 0$, $c_{2}=c_{3}=\frac{2}{3}$
$\begin{array}[]{c|ccc}0&\\ \frac{2}{3}&\frac{2}{3}\\ \frac{2}{3}&\frac{2}{3}-\frac{1}{4b_{3}}&\frac{1}{4b_{3}}\\ \hline\cr&\frac{1}{4}&\frac{3}{4}-b_{3}&b_{3}\\ \end{array}$
III. $b_{3}\neq 0$, $c_{2}=\frac{2}{3}$,$c_{3}=0$
$\begin{array}[]{c|ccc}0&\\ \frac{2}{3}&\frac{2}{3}\\ 0&-\frac{1}{4b_{3}}&\frac{1}{4b_{3}}\\ \hline\cr&\frac{1}{4}-b_{3}&\frac{3}{4}&b_{3}\\ \end{array}$
3 Explicit RBF Runge-Kutta Methods
In this section, we present two variants of Runge-Kutta methods with MQ and IMQ-radial basis functions wherein the shape parameter $\epsilon\in\mathbb{R}\cup i\mathbb{R}$. We will further introduce the mathematical expressions for the proposed methods, respective truncation error terms and the optimal value of the shape parameter.
3.1 Two-Stage Third-Order Methods
(a) MQ
The two-stage MQ-RBF Runge-Kutta method can be written as
$v_{n+1}=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}),$ (3.1)
with,
$k_{1}$ $=f(t_{n},v_{n}),$
$k_{2}$ $=f\left(t_{n}+c_{2}h,(v_{n}+hk_{1}a_{21})\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}\right),$
where $\epsilon_{n}^{2}$ shape parameter and $\epsilon_{n}^{2}=0$ reduces the method to the classical form which is a classical Runge-Kutta method. Another form of the above method is
[…]
By Taylor’s series expansion, the local truncation error becomes
$\tau$ $=f+\frac{h}{2}(f_{t}+ff_{u})+\frac{h^{2}}{3!}(f_{tt}+f_{tu}f+f(f_{tu}+ff_{uu})+f_{u}(f_{t}+ff_{u}))-b_{1}f$
$-b_{2}\left(f+c_{2}hf_{t}+\left(hfa_{21}-\frac{\epsilon_{n}^{2}a_{21}^{2}h^{2}u_{n}}{2}-\frac{\epsilon_{n}^{2}a_{21}^{2}h^{2}b_{2}c_{2}h^{3}}{2}\right)f_{u}\right)+\mathcal{O}(h^{3})$
$=\left(1-b_{1}-b_{2}\right)f+\left[\left(\frac{1}{2}-b_{2}c_{2}\right)f_{t}+\left(\frac{1}{2}-a_{21}b_{2}\right)f_{u}f\right]\,h$
$+\bigg{[}\left(\frac{1}{6}-\frac{1}{2}b_{2}c_{2}^{2}\right)f_{tt}+\left(\frac{1}{3}-a_{21}b_{2}c_{2}\right)f_{tu}f$
$+\left(\frac{1}{6}-\frac{1}{2}a_{21}^{2}b_{2}\right)f^{2}f_{uu}\bigg{]}h^{2}+\frac{(f_{t}+f_{u}f)f_{u}}{6}-\frac{\epsilon_{n}^{2}a_{21}^{2}b_{2}u_{n}f_{u}}{2}h^{2}+\mathcal{O}(h^{3}).$
The method will be third-order accurate if $a_{21}=c_{2}=\frac{2}{3}$, $b_{1}=\frac{1}{4}$, $b_{2}=\frac{3}{4}$ and
$\epsilon_{n}^{2}=\frac{f_{t}+ff_{u}}{u_{n}}.$
(b) IMQ
The two-stage IMQ-RBF Runge-Kutta method is of the form
$v_{n+1}=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}),$ (3.2)
with,
$k_{1}$ $=f(t_{n},v_{n}),$
$k_{2}$ $=f\left(t_{n}+c_{2}h,\sqrt{(1+\epsilon_{n}^{2}a_{21}^{2}h^{2})}hk_{1}+\frac{v_{n}}{\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}}\right),$
where $\epsilon_{n}^{2}$ is shape parameter and equating to zero leads the above method to the classical RK-2 method. This modified method can be expressed by the augmented Butcher tableau,
$\begin{array}[]{c|cc|c}0&&\\ c_{2}&a_{21}&&\epsilon_{n}^{2}\\ \hline\cr&b_{1}&b_{2}\end{array}$ (3.3)
where $\epsilon_{n}^{2}$, $a_{ij}$, $b_{i}$, and $c_{i}$ carry conventional definition. This two-stage method can be rewritten as
$v^{(1)}$ $=\sqrt{(1+\epsilon_{n}^{2}a_{21}^{2}h^{2})}hk_{1}+\frac{v_{n}}{\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}},$
$v_{n+1}$ $=v_{n}\left(1-\frac{b_{1}}{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}\right)+\frac{b_{1}v^{(1)}}{\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}}+b_{2}hf\left(t_{n}+c_{2}h,v^{(1)}\right).$
By Taylor’s series expansion, we have local truncation error as
$\tau$ $=\frac{u_{n+1}-u_{n}}{h}-(b_{1}k_{1}+b_{2}k_{2})$
$=f+\frac{h}{2}(f_{t}+ff_{u})+\frac{h^{2}}{3!}(f_{tt}+f_{tu}f+f(f_{tu}+ff_{uu})+f_{u}(f_{t}+ff_{u}))-b_{1}f$
$-b_{2}(f+c_{2}hf_{t}+\left(hfa_{21}-\frac{\epsilon_{n}^{2}a_{21}^{2}h^{2}u_{n}}{2}\right)+\mathcal{O}(h^{3})$
$=(1-b_{1}-b_{2})f+\left[\left(\frac{1}{2}-b_{2}c_{2}\right)f_{t}+\left(\frac{1}{2}-a_{21}b_{2}\right)f_{u}f\right]\,h$
$+\biggl{[}\left(\frac{1}{6}-\frac{1}{2}b_{2}c_{2}^{2}\right)f_{tt}+\left(\frac{1}{3}-a_{21}b_{2}c_{2}\right)f_{tu}f$
$+\left(\frac{1}{6}-\frac{1}{2}a_{21}^{2}b_{2}\right)f^{2}f_{uu}\biggl{]}h^{2}+(f_{t}+f_{u}f)f_{u}+\frac{\epsilon_{n}^{2}a_{21}^{2}b_{2}u_{n}f_{u}}{2}h^{2}+\mathcal{O}(h^{3}).$
The method will be third-order accurate if $a_{21}=c_{2}=\frac{2}{3}$, $b_{1}=\frac{1}{4}$, $b_{2}=\frac{3}{4}$ and
$\epsilon_{n}^{2}=\frac{f_{t}+ff_{u}}{u_{n}}.$
3.2 Three-stage forth-order methods
(a) MQ
$v_{n+1}=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}+b_{3}k_{3}),$ (3.4)
where,
$k_{1}$ $=f(t_{n},v_{n}),$
$k_{2}$ $=f\left(t_{n}+c_{2}h,(v_{n}+hk_{1}a_{21})\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\right),$
$k_{3}$ $=f\left(t_{n}+c_{3}h,(v_{n}+h(k_{1}a_{31}+k_{2}a_{32}))\sqrt{1+\epsilon_{n3}^{2}a_{21}^{2}h^{2}}\right).$
The method can be written as
$v^{(1)}$ $=\left(v_{n}+hfa_{21}\right)\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}},$
$v^{(2)}$ $=v_{n}\left(1-\frac{a_{31}}{a_{21}}\right)+\left[\frac{a_{31}v^{(1)}}{a_{21}\sqrt{1+\epsilon_{n}^{2}c_{2}^{2}h^{2}}}+ha_{32}f(t_{n}+c_{2}h,v^{(1)})\right]\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}},$
$v_{n+1}$ $=v_{n}\left[1-b_{1}h-\frac{b_{2}}{a_{32}h\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}+\frac{a_{31}}{a_{32}}\right]+v^{(1)}\left[\frac{b_{1}}{a_{21}\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}}-\frac{b_{2}a_{31}}{a_{32}\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}}\right]$
$\quad+\frac{v^{(2)}}{a_{32}h\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}+b_{3}f(t_{n}+c_{3}h,v^{(2)}).$
Using Taylor’s series expansion, the local truncation error derived as
$\tau_{n}$ $=\frac{u_{n+1}-u_{n}}{h}-(b_{1}k_{1}+b_{2}k_{2}+b_{3}k_{3})$
$=(1-b_{1}-b_{2}-b_{3})+h\left(-a_{21}b_{2}ff_{u}-a_{31}b_{3}ff_{u}-a_{32}b_{3}ff_{u}-b_{2}c_{2}f_{t}-b_{3}c_{3}f_{t}+\frac{ff_{u}}{2}+\frac{f_{t}}{2}\right)$
$\quad+h^{2}\Bigg{(}-\frac{a_{21}^{2}b_{2}f^{2}f_{uu}}{2}-a_{21}a_{32}b_{3}ff_{u}^{2}-a_{21}b_{2}c_{2}ff_{ut}-\frac{a_{31}^{2}b_{3}f^{2}f_{uu}}{2}-\frac{a_{31}a_{32}b_{3}f^{2}f_{uu}}{2}$
$\quad-a_{31}b_{3}c_{3}ff_{ut}-\frac{a_{32}^{2}b_{3}f^{2}f_{uu}}{2}-a_{32}b_{3}c_{2}f_{t}f_{u}-a_{32}b_{3}c_{3}ff_{ut}-\frac{b_{2}c_{2}^{2}}{2}v_{\epsilon 2}f_{u}-\frac{b_{2}c_{2}^{2}f_{tt}}{2}-\frac{b_{3}c_{3}^{2}v_{\epsilon 3}f_{u}}{2}$
$\quad-\frac{b_{3}c_{3}^{2}f_{tt}}{2}+\frac{f^{2}f_{uu}}{6}+\frac{ff_{u}^{2}}{6}+\frac{ff_{ut}}{3}+\frac{f_{t}f_{u}}{6}+\frac{f_{tt}}{6}\Bigg{)}$
$\quad+h^{3}\Bigg{(}-\frac{a_{21}^{2}b_{2}f^{2}f_{uu}}{2}-a_{21}a_{32}b_{3}ff_{u}^{2}-a_{21}b_{2}c_{2}ff_{ut}-\frac{a_{31}^{2}b_{3}f^{2}f_{uu}}{2}-\frac{a_{31}a_{32}b_{3}f^{2}f_{uu}}{2}$
$\quad-a_{31}b_{3}c_{3}ff_{ut}-\frac{a_{32}^{2}b_{3}f^{2}f_{uu}}{2}-a_{32}b_{3}c_{2}f_{t}f_{u}-a_{32}b_{3}c_{3}ff_{ut}-\frac{b_{2}c_{2}^{2}}{2}v_{\epsilon 2}f_{u}-\frac{b_{2}c_{2}^{2}f_{tt}}{2}-\frac{b_{3}c_{3}^{2}v_{\epsilon 3}f_{u}}{2}$
$\quad-\frac{b_{3}c_{3}^{2}f_{tt}}{2}+\frac{f^{2}f_{uu}}{6}+\frac{ff_{u}^{2}}{6}+\frac{ff_{ut}}{3}+\frac{f_{t}f_{u}}{6}+\frac{f_{tt}}{6}\Bigg{)}+\mathcal{O}(h^{3})$
As the conditions (2.2), (2.7) and
$b_{2}c_{2}^{2}\epsilon_{n2}^{2}+b_{3}c_{3}^{2}\epsilon_{n3}^{2}=0,$
are imposed the above local truncation error term vanishes. There are three cases where different parameter values makes the method forth order accurate.
  1. I.
    The coefficient of $h^{3}$ become zero when
    $a_{21}=c_{2}=\frac{1}{3},a_{31}=-\frac{5}{12},a_{32}=\frac{5}{4},c_{3}=\frac{5}{6},b_{1}=\frac{1}{10},b_{2}=\frac{1}{2},b_{3}=\frac{2}{5}.$
    $\epsilon^{2}_{n2}$ $=\frac{ff_{t}f_{uu}-3ff_{u}^{3}-ff_{u}f_{tu}-3f_{t}f_{u}^{2}+f_{t}f_{tu}-f_{tt}f_{u}}{vff_{uu}-2vf_{u}^{2}+vf_{tu}+ff_{u}},$
    $\epsilon^{2}_{n3}$ $=-\frac{1}{5}\epsilon^{2}_{n2}.$
  2. II.
    The coefficient of $h^{3}$ become zero when
    $a_{21}=c_{2}=1,a_{31}=\frac{1}{4},a_{32}=\frac{1}{4},c_{3}=\frac{1}{2},b_{1}=\frac{1}{6},b_{2}=\frac{1}{6},b_{3}=\frac{2}{3}.$
    $\epsilon^{2}_{n2}$ $=\frac{ff_{t}f_{uu}+ff_{u}^{3}-ff_{u}f_{tu}+f_{t}f_{u}^{2}+f_{t}f_{tu}-f_{tt}f_{u}}{vff_{uu}+2vf_{u}^{2}+vf_{tu}+ff_{u}},$
    $\epsilon^{2}_{n3}$ $=-\epsilon^{2}_{n2}.$
  3. III.
    The coefficient of $h^{3}$ become zero when
    $a_{21}=c_{2}=\frac{1}{2},a_{31}=0,a_{32}=c_{3}=\frac{3}{4},c_{3}=1,b_{1}=\frac{2}{9},b_{2}=\frac{1}{3},b_{3}=\frac{4}{9}.$
    $\epsilon^{2}_{n2}$ $=\frac{-\frac{1}{3}f^{3}f_{uuu}-f^{2}f_{tuu}-4ff_{u}^{3}-ff_{ttu}-4f_{t}f_{u}^{2}-\frac{1}{3}f_{ttt}}{vff_{uu}-4vf_{u}^{2}+vf_{tu}+ff_{u}}$
    $\epsilon^{2}_{n3}$ $=-\frac{1}{3}\epsilon^{2}_{n2}.$
(b) IMQ
$v_{n+1}=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}+b_{3}k_{3}),$ (3.5)
where,
$k_{1}$ $=f(t_{n},v_{n}),$
$k_{2}$ $=f\left(t_{n}+c_{2}h,\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}hf_{n}+\frac{v_{n}}{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}\right),$
$k_{3}$ $=f\left(t_{n}+c_{3}h,\sqrt{(1+\epsilon_{n3}^{2}c_{3}^{2}h^{2})}h(a_{31}k_{1}+a_{32}k_{2})+\frac{v_{n}}{\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}\right).$
Truncation error is
$\tau_{n}=$ $\frac{u_{n+1}-u_{n}}{h}-(b_{1}k_{1}+b_{2}k_{2}+b_{3}k_{3})$
$=(1-b_{1}-b_{2}-b_{3})f+h\left(-a_{21}b_{2}ff_{u}-a_{31}b_{3}ff_{u}-a_{32}b_{3}ff_{u}-b_{2}c_{2}f_{t}-b_{3}c_{3}f_{t}+\frac{ff_{v}}{2}+\frac{f_{t}}{2}\right)$
$+h^{2}\biggl{(}-\frac{a_{21}^{2}b_{2}f^{2}f_{uu}}{2}-a_{21}a_{32}b_{3}ff_{u}^{2}-a_{21}b_{2}c_{2}ff_{ut}-a_{31}a_{32}b_{3}f^{2}f_{uu}-a_{31}b_{3}c_{3}ff_{ut}-\frac{a_{32}^{2}b_{3}f^{2}f_{uu}}{2}$
$-a_{32}b_{3}c_{2}f_{t}f_{v}-a_{32}b_{3}c_{2}f_{t}f_{u}-a_{32}b_{3}c_{3}ff_{ut}+\frac{\epsilon_{n2}^{2}b_{2}c_{2}^{2}u_{n}f_{u}}{2}-\frac{b_{2}c_{2}^{2}f_{tt}}{2}+\frac{\epsilon_{n3}^{2}b_{3}c_{3}^{2}u_{n}}{2}-\frac{b_{3}c_{3}^{2}f_{tt}}{2}$
$+\frac{f^{2}f_{uu}}{6}+\frac{ff_{u}^{2}}{6}+\frac{ff_{ut}}{3}+\frac{f_{t}f_{u}}{6}+\frac{f_{tt}}{6}\biggl{)}+h^{3}\biggl{(}-\frac{-a_{21}^{3}b_{2}f^{3}f_{uuu}}{6}-\frac{a_{21}^{2}a_{32}b_{3}f^{2}f_{u}f_{uu}}{2}$
$-a_{21}b_{2}c_{2}f^{2}f_{uut}-a_{21}a_{31}a_{32}b_{3}f^{2}f_{u}f_{uu}-a_{21}a_{32}^{2}b_{3}f^{2}f_{u}f_{uu}-a_{21}a_{32}b_{3}c_{2}ff_{u}f_{ut}$
$-a_{21}a_{32}b_{3}c_{3}ff_{u}f_{ut}+\frac{a_{21}b_{2}c_{2}^{2}u_{n}\epsilon_{n2}^{2}ff_{uu}}{2}-\frac{a_{21}b_{2}c_{2}^{2}\epsilon_{n2}^{2}ff_{u}}{2}-a_{21}b_{2}c_{2}^{2}ff_{utt}-\frac{a_{31}^{3}b_{3}f^{3}f_{uuu}}{6}$
$-\frac{a_{31}^{2}a_{32}b_{3}f^{3}f_{uuu}}{2}-\frac{a_{31}^{2}b_{3}c_{3}f^{2}f_{uut}}{2}-\frac{a_{31}a_{32}^{2}b_{3}f^{3}f_{uuu}}{2}-a_{31}a_{32}b_{3}c_{2}ff_{t}f_{uu}-\frac{a_{31}b_{3}c_{3}^{2}u_{n}\epsilon_{3n}^{2}ff_{uu}}{2}$
$-\frac{a_{31}b_{3}c_{3}^{2}\epsilon_{3n}^{2}ff_{u}}{2}-\frac{a_{31}b_{3}c_{3}^{2}ff_{utt}}{2}-a_{32}^{2}b_{3}c_{2}ff_{t}f_{uu}-\frac{a_{32}^{2}b_{3}c_{3}f^{2}f_{uut}}{2}+\frac{a_{32}b_{3}c_{2}^{2}u_{n}\epsilon_{2n}^{2}f_{u}^{2}}{2}-\frac{a_{32}b_{3}c_{2}^{2}f_{tt}f_{u}}{2}$
$-a_{32}b_{3}c_{2}c_{3}f_{t}f_{u}t-\frac{a_{32}b_{3}c_{3}^{3}u_{n}\epsilon_{3n}^{2}ff_{uu}}{2}-\frac{a_{32}b_{3}c_{3}^{2}ff_{utt}}{2}+\frac{b_{2}c_{2}^{2}u_{n}\epsilon_{2}^{2}f_{ut}}{2}-\frac{b_{2}c_{2}^{3}f}{6}+\frac{b_{3}c_{3}^{3}u_{n}f_{ut}\epsilon_{3n}^{2}}{2}$
$-\frac{b_{3}c_{3}^{3}f_{ttt}}{6}+\frac{f^{3}f_{uuu}}{24}+\frac{f^{2}f_{u}f_{u}u}{6}+\frac{f^{2}f_{vt}}{8}+\frac{ff_{t}f_{u}u}{8}+\frac{ff_{u}^{3}}{24}+\frac{5ff_{u}f_{ut}}{24}+\frac{ff_{ut}}{8}$
$+\frac{f_{t}f_{u}^{2}}{24}+\frac{f_{t}f_{ut}}{8}+\frac{f_{u}f_{tt}}{24}+\frac{f_{ttt}}{24}\biggl{)}+\mathcal{O}(h^{4}).$
Applying (2.2), (2.7) and
$b_{2}c_{2}^{2}\epsilon_{n2}^{2}+b_{3}c_{3}^{2}\epsilon_{n3}^{2}=0$
to the above local truncation error term vanish. Here, we have four cases.
  1. I.
    The coefficient of $h^{3}$ become zero when,
    $a_{21}=c_{2}=\frac{1}{2},a_{31}=-1,a_{32}=2,c_{3}=1,b_{1}=\frac{1}{6},b_{2}=\frac{2}{3},b_{3}=\frac{1}{6},$
    $\epsilon^{2}_{n2}=\frac{-f^{2}f_{u}f_{uu}-ff_{t}f_{uu}+ff_{u}^{3}-ff_{u}f_{tu}+f_{t}f_{u}^{2}-f_{t}f_{tu}}{uff_{uu}-uf_{u}^{2}+uf_{tu}-ff_{u}},$
    $\epsilon^{2}_{n3}=-\epsilon^{2}_{n2}.$
  2. II.
    The coefficient of $h^{3}$ become zero when
    $a_{21}=c_{2}=\frac{1}{3},a_{31}=-\frac{5}{12},a_{32}=\frac{5}{4},c_{3}=\frac{5}{6},b_{1}=\frac{1}{10},b_{2}=\frac{1}{2},b_{3}=\frac{2}{5},$
    $\epsilon^{2}_{n2}$ $=\frac{-ff_{t}f_{uu}+3ff_{u}^{3}+ff_{u}f_{tu}+3f_{t}f_{u}^{2}-f_{t}f_{tu}+f_{tt}f_{u}}{uff_{uu}-2uf_{u}^{2}+uf_{tu}-ff_{u}},$
    $\epsilon^{2}_{n3}$ $=-\frac{1}{5}\epsilon^{2}_{n2}.$
  3. III.
    The coefficient of $h^{3}$ become zero when
    $a_{21}=c_{2}=1,a_{31}=\frac{1}{4},a_{32}=\frac{1}{4},c_{3}=\frac{1}{2},b_{1}=\frac{1}{6},b_{2}=\frac{1}{6},b_{3}=\frac{2}{3},$
    $\epsilon^{2}_{n2}$ $=\frac{-ff_{t}f_{uu}-ff_{u}^{3}+ff_{u}f_{tu}-f_{t}f_{u}^{2}-f_{t}f_{tu}+f_{tt}f_{u}}{uff_{uu}+2vf_{u}^{2}+uf_{tu}-ff_{u}},$
    $\epsilon^{2}_{n3}$ $=-\epsilon^{2}_{n2}.$
  4. IV.
    The coefficient of $h^{3}$ become zero when
    $a_{21}=c_{2}=\frac{1}{2},a_{31}=0,a_{32}=c_{3}=\frac{3}{4},c_{3}=1,b_{1}=\frac{2}{9},b_{2}=\frac{1}{3},b_{3}=\frac{4}{9},$
    $\epsilon^{2}_{n2}$ $=\frac{\frac{1}{3}f^{3}f_{uuu}+f^{2}f_{tuu}+4ff_{u}^{3}+ff_{ttu}+4f_{t}f_{u}^{2}+\frac{1}{3}f_{ttt}}{uff_{uu}-4uf_{u}^{2}+uf_{tu}-ff_{u}},$
    $\epsilon^{2}_{n3}$ $=-\frac{1}{3}\epsilon^{2}_{n2}.$
4 Convergence of RBF RUNGE-KUTTA Method
Numerical methods derived in previous section can be applied on IVP if there exists unique solution. IVP of the form (1.1) has a unique solution if $f(t,u)$ satisfies Lipschitz continuity with respect to second-coordinate that is $|f(t,u_{1})-f(t,u_{2})|\leq L|u_{1}-u_{2}|$, where $L$ is Lipschitz constant. Now, we prove the convergence of above proposed methods. Note that, we solve the IVP on an interval $[a,b]$ which is divided by uniform mesh length $h=\dfrac{b-a}{N}$ as $N+1$ points with mesh grid points $a=t_{0}<t_{1}<...<t_{N}=b$ where $t_{n}=a+nh,n=0,1,...,N.$ We also denote $u_{n}=u(t_{n})$ as exact solution and $v_{n}$ as numerical approximation for the proposed RBF Runge-Kutta schemes.
Theorem 4.1.
If $\epsilon_{n2}^{2}$ is bounded for all $n=0,1,\dots,N-1$, then the two-stage MQ-RBF Runge-Kutta method (3.1) converges provided the method satisfies (2.2), (2.3) and (2.5).
Proof.
Suppose, $E_{n+1}$ be the error between $u_{n+1}$ and $v_{n+1}$ for fix $t=t_{n}$. Now,
$k_{1}(w)=f(t_{n},w),$
$k_{2}(w)=f\left(t_{n}+c_{2}h,(w+hk_{1}a_{21})\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\right).$
$u_{n+1}$ can be expressed as
$u_{n+1}=u_{n}+h(b_{1}k_{1}(u_{n})+b_{2}k_{2}(u_{n}))+h\tau_{n},$
where $\tau_{n}$ is truncation error. Since, $f(t,u)$ satisfies Lipschitz continuity in $u$, so,
$|k_{1}(u_{n})-k_{1}(v_{n})|=|f(t_{n},u_{n})-f(t_{n},v_{n})|\leq L|u_{n}-v_{n}|,$
and
$|k_{2}(u_{n})-k_{2}(v_{n})|$ $=\biggl{|}f\left(t_{n},(u_{n}+ha_{21}k_{1}(u_{n}))\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\right)-f\left(t_{n},(v_{n}+ha_{21}k_{1}(v_{n})\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\right)\biggl{|},$
$\leq L\biggl{|}\left(u_{n}+ha_{21}k_{1}(u_{n})\right)\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}-\left(v_{n}+ha_{21}k_{1}(v_{n})\right)\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\biggl{|}$
$\leq L\biggl{|}(u_{n}-v_{n})\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+a_{21}hL(u_{n}-v_{n})\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\biggl{|}$
$\leq L\left(\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}+a_{21}hL\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\right)|u_{n}-v_{n}|,$
$L$ is Lipschitz constant.
$E_{n+1}$ $=|u_{n+1}-v_{n+1}|$
$=|(u_{n}-v_{n})+h[b_{1}(k_{1}(u_{n})-k_{1}(v_{n}))+b_{2}(k_{2}(u_{n})-k_{2}(v_{n}))]+h\tau_{n}|$
$\leq|u_{n}-v_{n}|+h[b_{1}|k_{1}(u_{n})-k_{1}(v_{n})|+b_{2}|k_{2}(u_{n})-k_{2}(v_{n})|]+h|\tau_{n}|$
$\leq\left(1+hLb_{1}+b_{2}(1+hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\right)|u_{n}-v_{n}|+h|\tau_{n}|$
$=\phi_{n}E_{n}+h|\tau_{n}|$
with
$\phi_{n}$ $=1+b_{1}hL+b_{2}(1+a_{21}hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}$
$\leq 1+hL+e^{hL}\mathcal{C}_{2}h$
$\leq e^{hL}+e^{hL}\mathcal{C}_{2}h$
$=e^{hL}(1+\mathcal{C}_{2}h)$
$=e^{h(L+\mathcal{C}_{2})}$
where,
$\mathcal{C}_{2}=\sup_{\begin{subarray}{c}0<h\leq b-a\\ n=0,\dots,N-1\end{subarray}}\frac{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}{h},$
Here, $c_{2}$ is constant and $\epsilon_{n2}^{2}$ is bounded as it depends on $u$ and $u$ is bounded ensured by Lipschitz continuity. Thus $E_{n}$ satisfies
$E_{n}\leq\sum_{j=0}^{n-1}\varphi_{j}E_{0}+h\sum_{j=1}^{n-1}\left(\sum_{m=j}^{n-1}\varphi_{m}\right)|\tau_{j-1}|+h|\tau_{n-1}|.$
Observing,
$\sum_{m=j}^{n-1}\varphi_{m}\leq\sum_{m=0}^{n-1}\varphi_{m}\leq e^{h(L+C_{2})n}\leq e^{h(L+C_{2})N}=e^{(L+C_{2})(b-a)}$
substituting, $E_{0}=0$ and
$\|\tau\|_{\infty}=\max_{n=0,\dots,N-1}|\tau_{n}|,$
it concludes that for every $n=0,1,\dots,N$,
$E_{n}\leq he^{(L+C_{2})T}\sum_{j=0}^{n-1}\|\tau\|_{\infty}=nhe^{(L+C_{2})(b-a)}\|\tau\|_{\infty}\leq(b-a)e^{(L+C_{2})(b-a)}\|\tau\|_{\infty}.$
and hence
$\lim_{\begin{subarray}{c}h\to 0\\ Nh=b-a\end{subarray}}E_{n}=0.$
So, $v_{n}$ converges to $u_{n}$. ∎
Theorem 4.2.
If $\epsilon_{n2}^{2}$ and $\epsilon_{n3}^{2}$ are bounded for all $n=0,1,\dots,N-1$. then the three-stage MQ-RBF Runge-Kutta method (3.4), converges provided the method satisfies (2.2), (2.3) and (2.7).
Proof.
Suppose, $E_{n+1}$ be the error between $u_{n+1}$ and $v_{n+1}$ for fix $t=t_{n}$. Now,
$k_{1}(w)=f(t_{n},w),$
$k_{2}(w)=f(t_{n}+c_{2}h,(w+hk_{1}a_{21})\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}),$
$k_{3}(w)=f(t_{n}+c_{3}h,(w+h(k_{1}a_{31}+k_{2}a_{32}))\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}).$
$u_{n+1}$ can be expressed as ,
$u_{n+1}=u_{n}+h(b_{1}k_{1}(u_{n})+b_{2}k_{2}(u_{n})+b_{3}k_{3}(u_{n}))+h\tau_{n},$
where $\tau_{n}$ is truncation error. Since, $f(t,u)$ satisfies Lipschitz continuity with respect to $u$, so,
$|k_{1}(u_{n})-k_{1}(v_{n})|=|f(t_{n},u_{n})-f(t_{n},v_{n})|\leq L|u_{n}-v_{n}|,$
$|k_{2}(u_{n})-k_{2}(v_{n})|\leq L\left(\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+a_{21}hL\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\right)\left|u_{n}-v_{n}\right|,$
$|k_{3}(u_{n})-k_{3}(v_{n})|$ $=\bigg{|}f\left(t_{n},\left(u_{n}+h(a_{31}k_{1}(u_{n})+a_{32}k_{2}(u_{n})\right)\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right)$
$-f\left(t_{n},(v_{n}+h(a_{31}k_{1}(v_{n})+a_{32}k_{2}(v_{n}))\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right)\bigg{|}$
$\leq L\left|(u_{n}-v_{n})+\left(ha_{31}L|u_{n}-v_{n}|+ha_{32}L(1+hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}|u_{n}-v_{n}|\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right)\right|$
$\leq L\left(1+\left(ha_{31}+ha_{32}(1+hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}L\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right)\right)|u_{n}-v_{n}|,$
$L$ is Lipschitz constant.
$E_{n+1}$ $=|u_{n+1}-v_{n+1}|$
$=|(u_{n}-v_{n})+h[b_{1}(k_{1}(u_{n})-k_{1}(v_{n}))+b_{2}(k_{2}(u_{n})-k_{2}(v_{n}))+b_{3}(k_{3}(u_{n})-k_{3}(v_{n}))]+h\tau_{n}|$
$\leq|u_{n}-v_{n}|+h\left[b_{1}|k_{1}(u_{n})-k_{1}(v_{n})|+b_{2}|k_{2}(u_{n})-k_{2}(v_{n})|+b_{3}|k_{3}(u_{n})-k_{3}(v_{n})|\right]+h|\tau_{n}|$
$\leq\left(1+hLb_{1}+b_{2}(1+a_{21}hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\right.$
$\quad+\left.b_{3}hL\left(1+h\left(a_{31}+ha_{32}(1+hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\right)\right)\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right)\left|u_{n}-v_{n}\right|+h|\tau_{n}|$
$=\phi_{n}E_{n}+h|\tau_{n}|$
Here,
$\phi_{n}$ $=1+b_{1}hL+b_{2}(1+hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+b_{3}hL\left(1+(hLa_{31}+hLa_{32}(1+hL))\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right)$
$\leq 1+hL+b_{2}hL(1+hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+b_{3}hL\left(1+hLa_{31}+hLa_{32}(1+hL)\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right)$
$\leq e^{hL}+b_{2}hLe^{hL}\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+b_{3}hL{e^{hL}+hLe^{hL}}\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}$
$\leq e^{hL}\biggl{(}1+b_{2}hL\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\biggl{)}+b_{3}hL(1+hL)\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}$
$\leq e^{hL}(1+\mathcal{C}_{3}h)$
$\leq e^{h(L+\mathcal{C}_{3})}$
with,
$\mathcal{C}_{3}=\sup_{\begin{subarray}{c}0<h\leq b-a\\ n=0,\dots,N-1\end{subarray}}\frac{b_{2}L\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+b_{3}L(1+hL)\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}{h},$
Here, $c_{3}$ is constant and $\epsilon_{n2}^{2}$ and $\epsilon_{n3}^{2}$ are bounded as it depends on $u$ and $u$ is bounded ensured by Lipschitz continuity. Therefore $E_{n}$ satisfies
$E_{n}\leq\sum_{j=0}^{n-1}\varphi_{j}E_{0}+h\sum_{j=1}^{n-1}\left(\sum_{m=j}^{n-1}\varphi_{m}\right)|\tau_{j-1}|+h|\tau_{n-1}|.$
Clearly,
$\sum_{m=j}^{n-1}\varphi_{m}\leq\sum_{m=0}^{n-1}\varphi_{m}\leq e^{h(L+C_{2})n}\leq e^{h(L+C_{2})N}=e^{(L+C_{2})(b-a)}.$
Upon considering $E_{0}=0$ and
$\|\tau\|_{\infty}=\max_{n=0,\dots,N-1}|\tau_{n}|,$
it can be concluded that for every $n=0,1,\dots,N$,
$E_{n}\leq he^{(L+C_{2})T}\sum_{j=0}^{n-1}\|\tau\|_{\infty}=nhe^{(L+C_{2})(b-a)}\|\tau\|_{\infty}\leq(b-a)e^{(L+C_{2})(b-a)}\|\tau\|_{\infty}$
and hence
$\lim_{\begin{subarray}{c}h\to 0\\ Nh=b-a\end{subarray}}E_{n}=0.$
Theorem 4.3.
If $\epsilon_{n2}^{2}$ is bounded for all $n=0,1,\dots,N-1$. Then the two-stage IMQ-RBF Runge-Kutta method (3.2) converges (2.2), provided it satisfies (2.3) and (2.5).
Proof.
Suppose, $E_{n+1}$ be the error between $u_{n+1}$ and $v_{n+1}$ for fix $t=t_{n}$. Let
$k_{1}(w)=f(t_{n},w),$
$k_{2}(w)=f\biggl{(}t_{n}+c_{2}h,\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\,\,hc_{2}k_{1}+\frac{w}{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}\biggl{)}.$
$u_{n+1}$ can be expressed as,
$u_{n+1}=u_{n}+h(b_{1}k_{1}(u_{n})+b_{2}k_{2}(u_{n}))+h\tau_{n},$
where $\tau_{n}$ be the trancation error. Since $f(t,u)$ satisfies Lipschitz continuity with respect to $u$,
$|k_{1}(u_{n})-k_{1}(v_{n})|=|f(t_{n},u_{n})-f(t_{n},v_{n})|\leq L|u_{n}-v_{n}|,$
and
$|k_{2}(u_{n})-k_{2}(v_{n})|$ $=\Biggl{|}f\left(t_{n},\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\,\,hc_{2}k_{1}(u_{n})+\frac{u_{n}}{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}\right)$
$-f\left(t_{n},\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\,\,hc_{2}k_{1}(v_{n})+\frac{v_{n}}{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}\right)\Biggl{|}$
$\leq L\Biggl{|}\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\,\,hc_{2}f(u_{n})+\frac{u_{n}}{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}-\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\,\,hc_{2}f(v_{n})$
$-\frac{v_{n}}{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}\Biggl{|}$
$\leq L\biggl{|}\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\,\,hc_{2}(f(u_{n})-f(v_{n}))+\frac{u_{n}-v_{n}}{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}\biggl{|}$
$\leq L\biggl{|}\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\,\,hc_{2}L+\frac{1}{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}\biggl{|}|u_{n}-v_{n}|,$
where $L$ be Lipschitz constant. Now,
$E_{n+1}$ $=|u_{n+1}-v_{n+1}|$
$=|(u_{n}-v_{n})+h[b_{1}(k_{1}(u_{n})-k_{1}(v_{n}))+b_{2}(k_{2}(u_{n})-k_{2}(v_{n}))]+h\tau_{n}|$
$\leq|u_{n}-v_{n}|+h[b_{1}|k_{1}(u_{n})-k_{1}(v_{n})|+b_{2}|k_{2}(u_{n})-k_{2}(v_{n})|]+h|\tau_{n}|$
$\leq|u_{n}-v_{n}|+hb_{1}L|u_{n}-v_{n}|+hb_{2}L\left|\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\,\,hc_{2}L+\frac{1}{\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}}\right||u_{n}-v_{n}|+h|\tau_{n}|$
$=\phi_{n}E_{n}+h|\tau_{n}|$
with the expression
$\phi_{n}$ $=1+b_{1}hL+hb_{2}L\left(\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\,\,Lhc_{2}+\frac{1}{\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}}\right)$
$\leq 1+b_{1}hL+hb_{2}L\left(\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\,\,Lhc_{2}+1\right)$
$=1+b_{1}hL+hb_{2}L\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\,\,Lha_{21}+hb_{2}L$
$=1+hb_{1}L+hb_{2}L\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}Lha_{21}+hL(1-b_{1})$
$=1+hL+hb_{2}L\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\,\,Lha_{21}$
where,
$\mathcal{C}_{2}=\sup_{\begin{subarray}{c}0<h\leq b-a\\ n=0,\dots,N-1\end{subarray}}\frac{hL^{2}b_{2}a_{21}\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}{e^{hL}}.$
Here, $c_{2}$ is constant and $\epsilon_{n2}^{2}$ is bounded as it depends on $u$ and $u$ is bounded ensured by Lipschitz continuity of $f$. Thus,
$\phi_{n}$ $\leq e^{hL}+\mathcal{C}_{2}he^{hL}$
$\leq(1+\mathcal{C}_{2}h)e^{hL}$
$\leq e^{h(L+\mathcal{C}_{2})}$
Hence $E_{n}$ satisfies,
$E_{n}\leq\sum_{j=0}^{n-1}\varphi_{j}E_{0}+h\sum_{j=1}^{n-1}\left(\sum_{m=j}^{n-1}\varphi_{m}\right)|\tau_{j-1}|+h|\tau_{n-1}|.$
Here,
$\sum_{m=j}^{n-1}\varphi_{m}\leq\sum_{m=0}^{n-1}\varphi_{m}\leq e^{h(L+C_{2})n}\leq e^{h(L+C_{2})N}=e^{(L+C_{2})(b-a)}.$
Assuming, $E_{0}=0$ and
$\|\tau\|_{\infty}=\max_{n=0,\dots,N-1}|\tau_{n}|,$
it follows that for every $n=0,1,\dots,N$,
$E_{n}\leq he^{(L+C_{2})T}\sum_{j=0}^{n-1}\|\tau\|_{\infty}=nhe^{(L+C_{2})(b-a)}\|\tau\|_{\infty}\leq(b-a)e^{(L+C_{2})(b-a)}\|\tau\|_{\infty}.$
Hence,
$\lim_{\begin{subarray}{c}h\to 0\\ Nh=b-a\end{subarray}}E_{n}=0.$
Theorem 4.4.
If $\epsilon_{n2}^{2}$ and $\epsilon_{n3}^{2}$ are bounded for all $n=0,1,\dots,N-1$. Then the three-stage RBF Runge-Kutta method (3.5), converges provided the method satisfies (2.2), (2.3) and (2.7).
Proof.
Suppose, $E_{n+1}$ be the error $E_{n+1}$ between $u_{n+1}$ and $v_{n+1}$ for fix $t=t_{n}$. Now,
$k_{1}(w)=f(t_{n},w),$
$k_{2}(w)=f\left(t_{n}+c_{2}h,\sqrt{(1+\epsilon_{n2}^{2}a_{21}^{2}h^{2})}hk_{1}+\frac{w}{\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}}\right),$
$k_{3}(w)=f\left(t_{n}+c_{3}h,\sqrt{(1+\epsilon_{n3}^{2}a_{31}^{2}h^{2})}hk_{1}+\frac{w}{\sqrt{1+\epsilon_{n3}^{2}c_{2}^{2}h^{2}}}\right).$
$u_{n+1}$ can be expressed as
$u_{n+1}=u_{n}+h(b_{1}k_{1}(u_{n})+b_{2}k_{2}(u_{n})+b_{3}k_{3}(u_{n}))+h\tau_{n},$
where, $\tau_{n}$ is trancation error. Since $f(t,u)$ is Lipschitz continuous in u , so
$|k_{1}(u_{n})-k_{1}(v_{n})|\leq L|u_{n}-v_{n}|,$
$|k_{2}(u_{n})-k_{2}(v_{n})|\leq L\left|\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}\,\,hk_{1}a_{21}+\frac{1}{\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}}\right||u_{n}-v_{n}|,$
$|k_{3}(u_{n})-k_{3}(v_{n})|$ $=\Bigg{|}f\left(t_{n}+c_{3}h,\ \sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\,\,h(a_{31}k_{1}(u_{n})+a_{32}k_{2}(u_{n}))+\frac{u_{n}}{\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}\right)$
$\quad-f\left(t_{n}+c_{3}h,\ \sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\,\,h(a_{31}k_{1}(v_{n})+a_{32}k_{2}(v_{n}))+\frac{v_{n}}{\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}\right)\Bigg{|}$
$\leq L\Bigg{|}\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}h\left(a_{31}(k_{1}(u_{n})-k_{1}(v_{n}))+a_{32}(k_{2}(u_{n})-k_{2}(v_{n}))\right)+\frac{u_{n}-v_{n}}{\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}\Bigg{|}$
$\leq L\left[\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\,\,ha_{31}L|u_{n}-v_{n}|+\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}\,\,ha_{32}L|u_{n}-v_{n}|\right.$
$\quad\left.+\frac{L|u_{n}-v_{n}|}{\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}}+\frac{|u_{n}-v_{n}|}{\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}\right].$
where, $L$ is Lipschitz constant. Now,
$E_{n+1}$ $=|u_{n+1}-v_{n+1}|$
$=|(u_{n}-v_{n})+h\left[b_{1}(k_{1}(u_{n})-k_{1}(v_{n}))+b_{2}(k_{2}(u_{n})-k_{2}(v_{n}))\right]+h\tau_{n}|$
$\leq|u_{n}-v_{n}|+hb_{1}L|u_{n}-v_{n}|+b_{2}hL\left(\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}\,\,hLa_{21}+\frac{1}{\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}}\right)|u_{n}-v_{n}|$
$+b_{3}hL\Biggr{[}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\,\,ha_{31}L+\sqrt{1+\epsilon_{n3}h^{2}c_{3}^{2}}\,\,ha_{32}\Biggr{(}L^{2}\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}\,\,ha_{21}+\frac{L}{\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}}\Biggl{)}\Biggr{]}|u_{n}-v_{n}|$
Now,
$\phi_{n}$ $=1+b_{1}hL+hb_{2}L\left(\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}\,\,hLa_{21}+\frac{1}{\sqrt{1+\epsilon_{n2}^{2}h^{2}a_{21}^{2}}}\right)$
$\quad+b_{3}hL\left[\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\,\,ha_{31}L+\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\,\,ha_{32}\left(L^{2}\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}\,\,ha_{21}+\frac{L}{\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}}\right)\right]$
$=1+hL\left(b_{1}+\frac{b_{2}}{\sqrt{1+\epsilon_{n2}^{2}h^{2}a_{21}^{2}}}+\frac{b_{3}}{\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}}\right)+h^{2}L^{2}\left(b_{2}a_{21}\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}+b_{3}a_{31}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\right.$
$\left.+a_{32}b_{3}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}\right)+b_{3}L^{3}h^{3}a_{21}a_{32}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\sqrt{1+\epsilon_{n2}^{2}h^{2}a_{21}^{2}}$
$\leq 1+hL+\frac{h^{2}L^{2}}{2}-\frac{h^{2}L^{2}}{2}\left(b_{2}a_{21}\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}+b_{3}a_{31}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\right.$
$\left.+a_{32}b_{3}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\right)+b_{3}L^{3}h^{3}a_{21}a_{32}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\sqrt{1+\epsilon_{n2}^{2}h^{2}a_{21}^{2}}$
$\leq(1+\mathcal{C}_{3}h)e^{hL}$
$\leq e^{h(1+\mathcal{C}_{3})}$
where,
$\mathcal{C}_{3}=\sup_{\begin{subarray}{c}0<h\leq b-a\\ n=0,\dots,N-1\end{subarray}}\frac{\begin{aligned} &h^{2}L^{3}b_{3}c_{2}a_{32}\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}-\frac{h^{2}L^{2}}{2}\biggl{(}b_{2}c_{2}\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\\ &+b_{3}a_{31}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}+b_{3}a_{32}\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\biggr{)}\end{aligned}}{e^{hL}},$
Here, $b_{3}$, $c_{2}$, $c_{3}$, $a_{32}$ is constant and $\epsilon_{n2}^{2}$ and $\epsilon_{n3}^{2}$ is bounded as it depends on $u$ and $u$ is bounded ensured by Lipschitz continuity. Therefore the term $E_{n}$ satisfies
$E_{n}\leq\sum_{j=0}^{n-1}\varphi_{j}E_{0}+h\sum_{j=1}^{n-1}\left(\sum_{m=j}^{n-1}\varphi_{m}\right)|\tau_{j-1}|+h|\tau_{n-1}|.$
Clearly,
$\sum_{m=j}^{n-1}\varphi_{m}\leq\sum_{m=0}^{n-1}\varphi_{m}\leq e^{h(L+C_{2})n}\leq e^{h(L+C_{2})N}=e^{(L+C_{2})(b-a)}.$
With $E_{0}=0$ and
$\|\tau\|_{\infty}=\max_{n=0,\dots,N-1}|\tau_{n}|,$
it can be concluded that for every $n=0,1,\dots,N$,
$E_{n}\leq he^{(L+C_{2})T}\sum_{j=0}^{n-1}\|\tau\|_{\infty}=nhe^{(L+C_{2})(b-a)}\|\tau\|_{\infty}\leq(b-a)e^{(L+C_{2})(b-a)}\|\tau\|_{\infty}.$
Hence,
$\lim_{\begin{subarray}{c}h\to 0\\ Nh=b-a\end{subarray}}E_{n}=0.$
5 Stability regions
Stability regions indicates to the set of values in complex plane where absolute value of the stability polynomial is bounded by one. $y^{\prime}=\lambda y$ is taken as the test function. Substituting the test function in the expression of the method and replacing $\lambda h$ by $z$, $h$ is step length, we obtain ratio of $v_{n+1}$ and $v_{n}$. The ratio is termed as stability polynomial and it is in the form $v_{n+1}=R(z)v_{n}$, where $R(z)$ is stability polynomial. Stability region is defined by $|R(z)|\leq 1$. Bounding the stability polynomial by one ensures that the error will not grow unboundedly. Stability polynomials of the two, three, and four stage claasical Runge-Kutta methods are as follows
$R(z)$ $=1+z+\frac{1}{2}z^{2},$ (5.1)
$R(z)$ $=1+z+\frac{1}{2}z^{2}+\frac{1}{6}z^{3},$ (5.2)
$R(z)$ $=1+z+\frac{1}{2}z^{2}+\frac{1}{6}z^{3}+\frac{1}{24}z^{4},$ (5.3)
which in general indicate the absolute stability region. The stability polynomial for two stage MQ RBF Runge-Kutta method (3.1) is given by
$R(z)=1+z+\frac{z^{2}}{2}+\frac{z^{3}}{6}+\frac{z^{4}}{9}.$
The stability polynomial for three stage MQ-RBF Runge-Kutta Method (3.4) are obtained as
  1. I.
    $R(z)=1+\frac{z}{10}+\frac{z}{2}\sqrt{1+\frac{z^{2}}{3}}+\frac{2}{5}\left[2z-\frac{5}{12}z^{2}+\frac{5z^{2}}{4}(1+\frac{z}{3}\sqrt{1+\frac{z^{2}}{3}}\right],$
  2. II.
    $R(z)=1+\frac{2z}{9}+\frac{z}{2}(1+\frac{z}{3})\sqrt{1+\frac{z^{2}}{3}}+\frac{2}{5}\left[2z+\frac{3}{4}(z^{2}+\frac{z^{2}}{3})\sqrt{1+\frac{z^{2}}{3}}\right],$
  3. III.
    $R(z)=1+\frac{z}{6}+\frac{1}{6}(z+z^{2})\sqrt{1+\frac{z^{2}}{3}}+\frac{4z}{3}+\frac{z^{2}}{6}+\frac{1}{6}(z+z^{2})\sqrt{1+\frac{z^{2}}{3}}.$
(a) Stability regions of RK2 (red) and IMQ-RK2 (blue).
(b) Stability regions of RK2 (red) and MQ-RK2 (blue).
(c) Stability regions of RK3 and IMQ-RK3.
(d) Stability regions of RK3 and MQ-RK3.
Figure 1: Stability region for various classical, MQ and IMQ-Runge Kutta schemes
The stability polynomial of the two stage IMQ RBF Runge-Kutta method (3.2) is
$R(z)=1+\frac{z}{4}+\frac{1}{2}z^{2}\sqrt{(1-\frac{4z^{2}}{9})}+\frac{3}{4}\frac{z}{\sqrt{(1-\frac{4z^{2}}{9}})}.$
The stability polynomial of the two-stage IMQ RBF Runge-Kutta method (3.5) are
  1. I.
    $R(z)=$ $1+\frac{z}{6}+\frac{2}{3}\left(z^{2}\sqrt{1-\frac{z^{2}}{4}}+\frac{z}{\sqrt{1-\frac{z^{2}}{4}}}\right)$
    $+\frac{1}{6}\left\{z\sqrt{1+z^{2}}\left(2z^{2}\sqrt{1-\frac{z^{2}}{4}}-z+\frac{2z}{\sqrt{1-\frac{z^{2}}{4}}}\right)+\frac{z}{\sqrt{1+z^{2}}}\right\},$
  2. II.
    $R(z)=$ $1+\frac{z}{10}+\frac{1}{2}\left(z^{2}\sqrt{1-\frac{z^{2}}{9}}+\frac{z}{\sqrt{1-\frac{z^{2}}{9}}}\right)$
    $+\frac{2}{5}\left\{z\sqrt{1+\frac{5}{36}z^{2}}\left(-\frac{5}{12}z+\frac{5z}{4}\left(z\sqrt{1-\frac{z^{2}}{9}}+\frac{1}{\sqrt{1-\frac{z^{2}}{9}}}\right)\right)+\frac{z}{\sqrt{1+\frac{5}{36}z^{2}}}\right\},$
  3. III.
    $R(z)=$ $1+\frac{z}{6}+\frac{1}{6}\left(z^{2}\sqrt{1-z^{2}}+\frac{1}{\sqrt{1-z^{2}}}\right)$
    $+\frac{2}{3}\left\{z\sqrt{1+\frac{z^{2}}{4}}\left(\frac{z}{4}+z\sqrt{1-z^{2}}+\frac{z}{\sqrt{1-z^{2}}}\right)+\frac{z}{\sqrt{1+\frac{z^{2}}{4}}}\right\},$
  4. IV.
    $R(z)=$ $1+\frac{2z}{9}+\frac{1}{3}\left(z^{2}\sqrt{1-\frac{z^{2}}{5}}+\frac{z}{\sqrt{1-\frac{z^{2}}{5}}}\right)$
    $+\frac{4}{9}\left\{z\sqrt{1-\frac{4z^{2}}{15}}\left(\frac{3z^{2}}{4}\sqrt{1-\frac{4z^{2}}{5}}+\frac{z}{\sqrt{1-\frac{z^{2}}{5}}}\right)+\frac{z}{\sqrt{1+\frac{4z^{2}}{15}}}\right\}.$
Figure1 indicates the stability regions of the proposed methods in comparison to classical RK methods. Figure1 (a) illustrates stability regions of two stage Runge-Kutta and IMQ-Runge Kutta method. Note that both the stability regions are in the negative real axis. Figure 1(b) shows stability region for RK-2 and MQ-RK-2 method. Figure1 (c) shows the stability region for various IMQ-RK-3 and RK-3 methods. The stability region for the MQ-RK-3 and RK-3 methods is shown in graph Figure 1(d). In each case, it is evident that the stability region for the classical Runge-Kutta method is larger than all enhanced Runge-Kutta methods.
6 Numerical results
In this section, numerical examples of MQ and IMQ RBF Runge-Kutta method are presented and discussed in support of the theory presented in previous sections. All calculations are performed with MATLAB R2022A. A comparison is followed with the corresponding classical Runge-Kutta method. All the errors preseneted here are calculated using $L_{1}-$norm.
Example 6.1
Consider the ordinary differential equation,
$\frac{du}{dt}=-u^{2},\quad 0<t\leq 1,$
with the initial condition;
$u(0)=1.$
The exact solution of the above ordinary differential equation is $u(t)=\frac{1}{t+1}$, decreasing in nature in $0<t\leq 1$. The global error evaluated at $t=1$ is reported for the standard RK2 method and its enhanced variants, namely MQ-RK2 and IMQ-RK2 in Table 1. Table 2 demonstrates that various RK3 methods achieve third-order accuracy, while different IMQ-RK3 methods attain fourth-order accuracy. Table 4 shows that the various MQ-RK3 methods described in Section 3 achieve fourth-order accuracy.
Table 1: Global error and order of accuracy by example 6.1 by RK2, MQ, IMQ-RK2.
N RK2 MQ-RK2 IMQ-RK2
Error Order Error Order Error Order
10 1.119140e-03 9.316803e-06 1.594597e-04
20 2.628612e-04 2.0900 1.487789e-06 2.6467 1.763600e-05 3.1766
40 6.368993e-05 2.0452 2.026835e-07 2.8759 2.074312e-06 3.0878
80 1.567527e-05 2.0226 2.626486e-08 2.9480 2.516187e-07 3.0433
160 3.888293e-06 2.0113 3.338011e-09 2.9761 3.098107e-08 3.0218
320 9.682818e-07 2.0056 4.205879e-10 2.9885 4.205879e-10 3.0109
Table 2: Global error and order of accuracy for example 6.1 by RK2 and IMQ-RK2.
N RK3 I IMQ-RK3 I RK3 II IMQ-RK3 II
Error Order Error Order Error Order Error Order
10 2.642520e-05 4.633848e-06 4.783447e-05 1.592061e-06
20 2.916366e-06 3.1797 2.573850e-07 4.1692 5.580618e-06 3.0999 9.390044e-08 4.0836
40 3.439349e-07 3.0840 1.509396e-08 4.0927 6.738067e-07 3.0246 5.681100e-09 4.0469
80 4.181551e-08 3.0400 9.130775e-10 4.0235 8.280145e-08 3.0152 3.492980e-10 4.0236
160 5.155666e-09 3.0185 5.614342e-11 4.0235 1.026114e-08 3.0125 2.165412e-11 4.0117
320 6.400902e-10 3.0098 3.480549e-12 4.0117 1.277089e-09 3.0063 1.347748e-12 4.0063
N RK3 III IMQ-RK3 III RK3 IV IMQ-RK3 IV
Error Order Error Order Error Order Error Order
10 4.716981e-05 5.617946e-06 4.786103e-05 2.274155e-06
20 5.548240e-06 3.0878 3.166580e-07 4.1490 5.583189e-06 3.0997 1.311771e-07 4.1157
40 6.720259e-07 3.0454 1.870708e-08 4.0813 6.739872e-07 3.0503 7.847559e-09 4.0631
80 8.270012e-08 3.0226 1.136403e-09 4.0140 8.281589e-08 3.0247 4.798086e-10 4.0317
160 1.025485e-08 3.0116 7.002599e-11 4.0204 1.026189e-08 3.0166 2.966272e-11 4.0157
320 1.276703e-09 3.0058 4.344858e-12 4.0105 1.277132e-09 3.0063 1.844919e-12 4.0076
Table 3: Global error and order of accuracy for example 6.1 by various MQ-RK3 methods.
N MQ-RK3 I MQ-RK3 II MQ-RK3 III
Error Order Error Order Error Order
10 4.536178e-06 1.609152e-06 3.020182e-06
20 2.720225e-07 4.0597 9.790250e-08 4.0388 1.799044e-07 4.0693
40 1.655922e-08 4.0380 6.007887e-09 4.0264 1.092505e-08 4.0415
80 1.020717e-09 4.0200 3.718700e-10 4.0140 6.727541e-10 4.0214
160 6.334999e-11 4.0101 2.312794e-11 4.0071 4.173539e-11 4.0107
320 3.945511e-12 4.0051 1.441736e-12 4.0038 2.599143e-12 4.0052
Example 6.2
Next, we consider the ordinary differential equation as
$\frac{du}{dt}=-4t^{3}u^{2},\quad-10<t\leq 0,\quad u(-10)=\frac{1}{10001}$
This is a stiff problem with the exact solution $u(t)=\frac{1}{t^{4}+1}$ where the solution exhibits rapid variation over the interval $[-10,0]$. The global error is evaluated at $t=0$. Table 4 shows that the numerical solution achieves second-order accuracy with the standard RK2 method and third-order accuracy with the enhanced variants MQ-RK2 and IMQ-RK2. As shown in Table 5, the various RK3 and IMQ-RK3 methods yield third and fourth order accuracy, respectively. These RK3 and IMQ-RK3 methods are discussed in Sections 2 and 3, respectively. Table 6 presents the global error and confirms fourth-order accuracy for the ODE using various MQ-RBF methods described in Section 3.
Table 4: Global error and order of accuracy for example 6.2 by RK2, MQ-RK2 and IMQ-RK2
N RK2 MQ-RK2 IMQ-RK2
Error Order Error Order Error Order
200 9.422354e-02 3.148990e-02 9.422354e-02
400 1.344393e-02 2.0900 4.055329e-03 2.9570 1.344393e-02 2.8091
800 1.738854e-03 2.0452 5.197901e-04 2.9638 1.738854e-03 2.9507
1600 2.201011e-04 2.0226 6.588221e-05 2.9800 2.201011e-04 2.9819
3200 2.766968e-05 2.0113 8.294091e-06 2.9897 2.766968e-05 2.9918
6400 3.468332e-06 2.0056 1.040483e-06 2.9948 3.468332e-06 2.9960
Table 5: Global error and order of accuracy for example 6.2 by RK2, IMQ-RK2.
N RK3 I IMQ-RK3 I RK3 II IMQ-RK3 II
Error Order Error Order Error Order Error Order
200 4.341821e-02 1.599661e-03 4.344246e-02 4.063955e-04
400 5.854329e-03 2.8907 1.047616e-04 3.9326 5.826218e-03 2.8985 2.661830e-05 3.9324
800 7.493172e-04 2.9659 6.705566e-06 3.9656 7.433671e-04 2.9704 1.703303e-06 3.9660
1600 9.460367e-05 2.9856 4.241174e-07 3.9828 9.369507e-05 2.9880 1.077257e-07 3.9829
3200 1.188195e-05 2.9931 2.666491e-08 3.9914 1.175767e-05 2.9944 6.777914e-09 3.9904
6400 1.488713e-06 2.9966 1.690263e-09 3.9796 1.472522e-06 2.9972 4.157718e-10 4.0270
N RK3 III IMQ-RK3 III RK3 IV IMQ-RK3 IV
Error Order Error Order Error Order Error Order
200 6.749350e-02 2.373662e-03 4.768849e-02 7.207373e-04
400 9.267981e-03 2.8644 1.666272e-04 3.8324 6.425003e-03 2.8919 4.757574e-05 3.9212
800 1.186675e-03 2.9653 1.068826e-05 3.9625 8.204686e-04 2.9692 3.055810e-06 3.9606
1600 1.496573e-04 2.9872 4.660953e-07 4.5193 1.034367e-04 2.9877 1.936161e-07 3.9803
3200 1.878283e-05 2.9942 2.926613e-08 3.9933 1.298126e-05 2.9942 1.219565e-08 3.9888
6400 2.352475e-06 2.9972 1.852862e-09 3.9814 1.625848e-06 2.9972 7.586857e-10 4.0067
Table 6: Global error and order of accuracy for example 6.2 by various MQ-RK3 methods.
N MQ-RK3 I MQ-RK3 II MQ-RK3 III
Error Order Error Order Error Order
200 4.647968e-04 4.766731e-04 1.093432e-04
400 2.913064e-05 3.9960 3.310703e-05 3.8478 5.926462e-06 4.2055
800 1.821534e-06 3.9993 2.212692e-06 3.9033 3.397134e-07 4.1248
1600 1.138538e-07 3.9999 1.411230e-07 3.9708 2.024166e-08 4.0689
3200 7.122291e-09 3.9987 8.862087e-09 3.9932 1.230801e-09 4.0397
6400 4.345316e-10 4.0348 5.673884e-10 3.9652 8.104606e-11 3.9247
Example 6.3
Consider,
$\frac{du}{dt}=\frac{2t^{2}-u}{t^{2}u-t},\quad 1<t\leq 2,\quad u(1)=2,$
The non-separable ordinary differential equation has the exact solution $u(t)=\frac{1}{t}+\sqrt{\frac{1}{t}^{2}+4t-4},$ which is increasing over the interval $t=1$ to $t=2$. The global error at the final time $t=2$ is computed using RK2, MQ-RK2, and IMQ-RK2 methods, as shown in Table 7. The order of accuracy is also calculated for these methods. RK2 achieves second-order accuracy, while MQ-RK2 and IMQ-RK2 attain third-order accuracy. Table 8 presents the global errors for various RK3 and IMQ-RK3 methods. The different IMQ-RK3 variants discussed in Section 3 outperform the standard RK3 methods in terms of both global error and accuracy. Table 9 shows the solution of the above ODE using various MQ-RK3 methods, demonstrating fourth-order accuracy.
Table 7: Global error and order of accuracy for example 6.3 by RK2, MQ-RK2 and IMQ-RK2.
N RK2 MQ-RK2 IMQ-RK2
Error Order Error Order Error Order
10 8.789865e-04 2.184352e-04 2.106559e-04
20 2.076179e-04 2.0819 2.542775e-05 3.1027 2.386215e-05 3.1421
40 5.048556e-05 2.0400 3.064144e-06 3.0528 2.836513e-06 3.0725
80 1.244991e-05 2.0197 3.757647e-07 3.0276 3.460363e-07 3.0351
160 3.091751e-06 2.0096 4.652147e-08 3.0139 4.272868e-08 3.0176
320 7.703476e-07 2.0048 5.787081e-09 3.0070 5.308747e-09 3.0088
Table 8: Global error and order of accuracy for example 6.3 by RK and IMQ-RK3.
N RK3 I IMQ-RK3 I RK3 II IMQ-RK3 II
Error Order Error Order Error Order Error Order
10 6.819460e-06 9.102354e-07 5.447556e-05 1.036981e-05
20 1.559008e-06 2.1290 7.556264e-08 3.5905 6.523486e-06 3.0619 1.007890e-06 3.3630
40 2.280679e-07 2.7731 5.461295e-09 3.7904 7.935154e-07 3.0393 6.019957e-08 4.0654
80 3.019340e-08 2.9172 3.659935e-10 3.8994 9.775898e-08 3.0210 7.278295e-08 -0.2738
160 3.869651e-09 2.9640 2.366685e-11 3.9509 1.212638e-08 3.0111 2.518786e-09 4.8528
320 4.891940e-10 2.9837 1.504130e-12 3.9759 1.509893e-09 3.0056 1.009757e-10 4.6406
N RK3 III IMQ-RK3 III RK3 IV IMQ-RK3 IV
Error Order Error Order Error Order Error Order
10 6.916137e-05 3.238976e-05 7.024289e-05 9.102354e-07
20 8.770902e-06 2.9792 2.286074e-06 3.8246 8.622116e-06 3.0262 7.556264e-08 3.5905
40 1.099991e-06 2.9952 1.427746e-07 4.0011 1.063681e-06 3.0190 5.461295e-09 3.7904
80 1.376718e-07 2.9982 8.792510e-09 4.0213 1.320313e-07 3.0101 3.659935e-10 3.8994
160 1.721441e-08 2.9995 5.430785e-10 4.0170 1.644235e-08 3.0054 2.366685e-11 3.9509
320 2.152050e-09 2.9998 3.370415e-11 4.0102 2.051421e-09 3.0027 1.504130e-12 3.9759
Table 9: Global error and order of accuracy for example 6.3 by various MQ-RK3 methods.
N MQ-RK3 I MQ-RK3 II MQ-RK3 III
Error Order Error Order Error Order
10 8.158208e-06 1.795212e-04 2.711265e-05
20 4.606486e-07 4.1465 4.752717e-06 5.2393 1.480904e-06 4.1944
40 2.749617e-08 4.0664 1.419743e-07 5.0651 8.654651e-08 4.0969
80 1.681109e-09 4.0317 5.324079e-09 4.7370 5.230366e-09 4.0485
160 1.039404e-10 4.0156 2.414269e-10 4.4629 3.214189e-10 4.0244
320 6.463274e-12 4.0073 1.257616e-11 4.2628 1.991918e-11 4.0122
7 Conclusions
In this study, we introduced MQ-RBF and IMQ-RBF-based Runge-Kutta methods, which demonstrate improved accuracy orders compared to classical Runge-Kutta schemes. A detailed analysis of convergence behavior, stability regions, and local convergence order was conducted. The results indicate that although the IMQ-RBF method incurs higher computational complexity, it consistently outperforms the MQ-RBF method in terms of both numerical error and convergence order. However, implementing fourth-stage, fifth-order Runge-Kutta methods with MQ and IMQ radial basis functions (RBFs) on standard CPU architectures presents significant memory challenges, complicating the search for optimal shape parameters. To address these limitations, the development of memory-efficient MQ and IMQ RBF-based fourth-order, five-stage Runge-Kutta methods is currently underway and will be presented in future work.